You are taking a 6-hour train ride. During the 1st part of the trip, the train moves at an average speed 65 miles per hour. During the 2nd part of the trip, the train moves an average speed of 40 miles per hour. Let L be the time it takes for the train to travel the 1st part of the trip. Write and simplify and expression in terms of t for the total distance you traveled in the train. Then find the total distance you traveled if the first part of the trip took 4 hours.
Answer by mananth(16949)

Total time = 6 hours
L = Time first part of trip
speed I part = 65 mph
Distance I part = 65 L
Time II part = 6 -L
Speed II part = 40 mph

Distance II part = 40(6-L)
Total Distance = 65L + (6-L)*40
=65L+240-40L
=25L+240
Total Distance =25L+240
If L=4
D= 100+240 =340 miles